I am not getting this at all. Here is the equation. One number is six more than another. Their sum is ten. Find both numbers.
I tried it the way Theo 5775 and it was so consing I think what I am need the formula without all the extra wording I want to formulate just plug in numbers

Answer by mananth(16946)

if one number is x
the other is 6 more than x
the numbers are x & (x+6)
sum of them is 10
x+x+6 =10
2x+6 =10
2x = 10-6
2x=4
/2
x=2
if x = 2 then the other is 8